OpenStudy (andertj21):

Which of the following is an arithmetic sequence? -2, 4, -6, 8, ... -8, -6, -4, -2, ... 2, 4, 8, 16, ...

OpenStudy (ikram002p):

what do u think ?

OpenStudy (andertj21):

c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

ur right it is c

OpenStudy (andertj21):

Thnx bby:) <3

OpenStudy (anonymous):

no problem... :)

OpenStudy (ikram002p):

nope , arithmetic sequence? must have constant common difference C is wrong answer since 16-8=8 8-4=4 4-2=3